OpenStudy (anonymous):

How do I solve this quadratic equation -4x^2=0

OpenStudy (anonymous):

x^2 = 0/-4 x^2 = 0 so x = 0

jhonyy9 (jhonyy9):

-4x^2 =0 /(-4) x^2 =0 x_1 = x_2 = 0